# WATonoBus: An All Weather Autonomous Shuttle

### v1

@@ III-C Decision Making
The objective of decision making module is to make safe driving behavior decisions based on the information from perception, prediction, localization modules, and HD map [29]. The high-level decisions will then serve as the inputs to the downstream modules to perform path planning, pullover/merging, and vehicle control. Within the service environment of WATonoBus, the driving behavior decisions are limited to Go/Stop decisions during normal driving and intersection handling, and additional Pullover/Merge, door open/close decisions at bus stop regions. Apart from normal driving decisions, the decision making module will also output emergency stop decisions as the safety fallback strategy. Furthermore, although all the interested obstacles will be reviewed through the decision making module, the final decision will be made with respect to the highest risk obstacle. To achieve the aforementioned objective, finite state machine (FSM) was chosen as the backbone in WATonoBus decision making module (See Fig. 8) because its strengths in structure modularity, interpretation transparency, and robustness to errors or unexpected events [30].
@@
The three important FSM components of the WATonoBus decision making module are states, state transition conditions, and decision-making logic in each state. In an FSM, a state represents a unique situation in the system. In this case, the states are defined based on the static road features in the bus service environment. Specifically, there are normal road segments, all-way stop T-shape intersections, and bus stop regions along the campus ring road. Therefore, the five major states in FSM are defined, namely normal driving, intersection handling, pullover, merging, and passenger on-boarding states. Moreover, to handle the unexpected events, an emergency state is added. A transition in the FSM is a change from one state to another. For WATonoBus decision making module, the transitions are mainly governed by the ego vehicle location and the perception results. Specifically, the ego vehicle location is used to compute the remaining distance to the coming static road features such as all-way stops or bus stop regions. For example, the FSM will change to the pullover state if the ego vehicle is approaching to a bus stop region and then a pullover path will be required. In addition, perception results contain the obstacles around the ego vehicle, indicating whether it is safe to resume bus normal driving. For instance, if there is no vehicle approaching from the rear, the FSM will switch from pullover state to merging state after waiting for passenger on-boarding. Lastly, decisions made under each state can base on different logic due to the traffic rules and vehicle control requirements. The traffic rules in each FSM state list the expected ego vehicle behaviors. To handle an intersection properly, a few traffic rules need to be satisfied in sequence. Assuming there is no pedestrian, the ego vehicle needs to first check if a stop is required for the intersection-waiting vehicle in its lane. It will then be required to stop at the intersection for a while till the intersection is clear. Depending on the arrival order, a decision of whether to yield for the waiting vehicle at other intersection entrances needs to be made. After clearing all these traffic rules check, the ego vehicle can finally start moving and let FSM switch from intersection handling to normal driving state. In addition to traffic rules, the decision-making logic in each state differ from the required vehicle control parameters. For instance, the maximum speed of merging is slower than that of normal driving to ensure the bus to carefully merge out of the bus region while be able to stop immediately for any vehicles suddenly come from the rear.
@@ III-D1 Global path planning
The global planner determines a high-level path from the current vehicle location to the destination, utilizing HD map information. This algorithm ensures a safe and efficient path, considering any vehicle or environmental constraints. As WATonoBus operates on a one-lane road, the global planner defines the right curb and center line as the right and left boundaries of the drivable space. The middle of this space is assigned as the reference path for the bus.
@@ III-D2 Local path planning
The task of local path planning is to come up with a more detailed path that follows the global path while taking into account the current state of the vehicle and any obstacles in its immediate surroundings. Moreover, this involves devising a path for the WATonoBus to safely execute pullover and merging maneuvers as needed. These are typically done using potential field [31] and Bezier curves [32, 33, 34]. With the potential field method, we have a mathematical framework to adjust the reference path of the bus whenever obstacles are detected within its drivable space. Following this, smooth segments are generated using Bezier curves from the modified waypoints obtained through the potential field method. For the pullover and merging maneuver, the waypoints are produced using the Bezier curves and some control points. By combining the potential field with Bezier curves, the WATonoBus can quickly and safely navigate through its environment and do its missions.
Following the local path, the objective of WATonoBus motion controller is to select a desired speed and steering angle for the system to track at each sampling instant that enables the vehicle to follow the reference path while safely interacting with obstacles. The longitudinal controller utilizes a feed-forward algorithm that considers the position and velocity of the ego vehicle relative to obstacles. Firstly, the position $s$ of the WATonoBus and other obstacles along the path is estimated. The distance difference $\Delta s$ between the bus and obstacles along the path is then calculated. If there are no obstacles in the vicinity of the bus, the desired velocity selected by the controller is updated incrementally at each sampling instant using
@@ III-E1 Safety Testing and Validation
During the offline development and testing stage, we simulated the parallel virtual environment of the campus ring road to validate our software across key scenarios. Safety validation in virtual traffic minimizes the risk of discovering high-impact software deficiencies in the real world. To encompass various challenging cases, we developed a scenario generator to explore potentially dangerous movements of vehicles and pedestrians. It considers various configuration parameters like the number of traffic participants, their motion and velocity profiles. We explore different combinations of these configuration parameters in testing scenario generation and then examine the performance of decision making and control modules for any potential failures. Any collisions and rule violations in the simulation will be reported and examined for further improvements.
III-E2 Safety Monitor and E-Stop
During real-time operation, system safety is maintained through reliability monitoring and fail-safe modules. Reliability monitoring module checks location, weather, and sensor performance, generating malfunction signals within the predefined Operational Design Domain (ODD) [39]. These signals indicate the current estimation of system reliability in perception, localization, and planning. The fail-safe module responds to warnings by prompting the operator to take over or initiating an emergency stop based on warning priority. Fig. 8 illustrates that all normal states can transition to an emergency stop through a software-level trigger. In the event of software freeze, an additional hardware trigger ensures a hard emergency stop.
III-E3 Black Box Recording
Our blackbox recording module is designed to enhance system learning and ensure continuous improvement through real world edge case recordings. This module maintains a rolling log of the last 30 seconds of operation, capturing crucial data on the shuttle’s state, sensor readings, and decision-making processes. A key feature of this module is its ability to recognize when a human operator takes over the controls. Once this event is detected, this module immediately stores the preceding 30 seconds of data into a rosbag, a file format in ROS used for storing ROS message data. This captured data provides valuable insights into the circumstances that led to the operator’s intervention. Simultaneously, the module emits a beeping sound to alert the operator of the recording process and the need to provide feedback. This is a crucial step in our process, as the operator is then given 30 seconds to provide a verbal summary of the situation that prompted the takeover. This feedback serves as a first-hand account of the event, aiding in our understanding of any gaps in the system’s autonomous decision-making capabilities and allowing us to refine and improve our technology further.
@@ IV Learning from Daily Operation
This section presents the handling procedure and lessons learned from both regular bus service and severe weather encountered in WATonoBus daily operations, followed by discussing the insights gained from these case studies.
@@ IV-A Bus Stop Handling
The passenger on/off boarding module ensures a comfortable bus service experience. After pullover at the bus stop, the shuttle automatically opens the door and uses the in-door sensor to detect potential passengers attempting to board or alight. After a continuous check and waiting for a sufficient time, the shuttle locks the door and transitions to the merging status.
The merging handling module combines rear radar and camera results for a comprehensive view of the traffic behind the ego vehicle, improving hazard detection and reaction capabilities. The radar provides reliable long-range detection, crucial for monitoring approaching traffic. It not only detects vehicles but also estimates their velocities, aiding in understanding their movement patterns. The rear camera complements these capabilities, serving as the final vote to prevent radar false alarms during merging. This fused data informs the vehicle’s DM process, calculating the optimal timing and speed for a smooth merge into traffic. Anticipating potential conflicts by considering the speed and trajectory of approaching vehicles, our system adjusts its strategy for safety, ensuring the well-being of shuttle bus passengers and other road users.
The pullover/merging path planning module ensures a smooth path for the ego vehicle during pullover at bus stops or merging maneuvers. Specific waypoints are crucial during pullover, determined considering the bus’s current position, heading, and the identified parking location. Bezier curves are utilized to create a smooth path between these waypoints. Similarly, for merging, waypoints are determined based on the bus’s starting position, orientation, and the target location along its reference path. These waypoints guide the creation of a smooth path using Bezier curves.
(a) Spot identification
(b) Pullover path planning
(c) Passenger handling
@@ IV-B Intersection Handling
Non-signalized intersections pose challenges due to dynamic traffic flow, right-of-way determination, and potential conflicts (See Fig. 11). Unlike signalized intersections, they lack explicit instructions, demanding advanced perception, DM, and control capabilities for autonomous shuttle buses. Non-signalized intersections rely on right-of-way rules, which can be influenced by factors such as vehicle priority, the order of arrival. Our WATonoBus records the order of arrival for vehicle actors to decide the road priority. When comparing with pedestrian and cyclist at crossings, the shuttle consider itself lowest road priority. Moreover, non-signalized intersections introduce the possibility of conflicts arising from simultaneous movements of different road users. Conflicts can occur due to misjudgment of gaps, ambiguous driver intentions, or violations of right-of-way. The autonomous shuttle bus must be equipped with robust perception systems and prediction modules to detect and track vehicles, pedestrians, and cyclists. To reach as safe and robust behavioral planning for the autonomous driving shuttle bus, we use event-driven technique that react the most potential conflict that guarantee the vehicle is always at safe states.
@@
Figure 11: Intersection handling
@@ IV-C Perception Performance under Snow
In autonomous vehicle systems, reliable object detection under adverse weather conditions remains a complex and critical challenge. Our WATonoBus perception system overcomes this challenge through the fusion of radar, LiDAR, and camera data, particularly in snowy settings. This integrated approach successfully detects objects in three disparate yet commonplace road scenarios: a van entering the roadway, a car in the opposite lane, and a pedestrian crossing a street, as shown in Fig. 12. For detecting a van entering the road, the radar’s capacity to measure distance and speed plays a crucial role, especially when fused with the camera’s high-definition imaging. The camera contributes valuable details like color and texture, allowing for a highly integrated and predictive object detection framework. This multi-sensory combination anticipates the van’s potential actions, enabling dynamic adjustments for safer autonomous navigation. In the case of a car in the opposite lane, snowy conditions can introduce additional complexities like fog, glare, and reduced contrast. Here, radar’s strong penetration capabilities prove indispensable for initial object detection and range estimation. Combined with the LiDAR’s noisy but detailed spatial data and the camera’s visual cues, the system doesn’t just detect the car; it precisely locates and tracks it, adding another layer of reliability and safety in treacherous weather conditions. Regarding a pedestrian on a crosswalk, snowy conditions often introduce noise into the LiDAR point clouds, making it more challenging to distinguish the shape and position of a pedestrian. In this setting, radar’s Doppler velocity data becomes invaluable. Radar’s robustness to weather interference complements the high-resolution LiDAR data, filtering the noise and ensuring a reliable and comprehensive multi-modal perception of the pedestrian. This fusion mitigates the limitations posed by snowy conditions, providing an enhanced layer of safety and accuracy.
@@
IV-D Learning and Discussion
The learning from the WATonoBus operational scenarios highlights the robustness and adaptability in addressing complex bus service tasks and adverse weather conditions. The pullover spot identification and merging handling modules showcase the system’s capability to navigate intricate tasks, ensuring safe and precise operations at bus stops. The integration of passenger on/off boarding mechanisms not only enhances user experience but also prioritizes safety, demonstrating a holistic approach to autonomous bus service. Moreover, the intelligent decision-making at non-signalized intersections reflects WATonoBus can dynamically determine right-of-way in complex traffic scenarios. The perceptual performance under snow adds another layer of strength to the WATonoBus design. The successful fusion of radar, LiDAR, and camera data in challenging weather conditions exemplifies the system’s reliance on multi-sensory inputs. Understanding the distinct role of each sensor in different scenarios emphasizes the system’s adaptability, ensuring reliable object detection and tracking under adverse weather. In essence, WATonoBus stands out as a resilient and intelligent autonomous bus service, capable of handling complex tasks with precision and adapting seamlessly to adverse weather conditions. The integration of advanced modules, coupled with a multi-sensory perception system, positions WATonoBus as a robust solution for safe and reliable autonomous transportation services.
Acknowledgment
The authors would like to acknowledge the financial support of Natural Sciences and Engineering Research Council of Canada (NSERC) and Canadian Foundation for Innovation (CFI) in this work.

### v2

@@ III-C Decision Making
The objective of the decision making module is to make safe driving decisions based on the information from perception, prediction, localization modules, and HD map [27, 28, 29]. The high-level decisions will then serve as the inputs to the downstream modules to perform path planning, pullover/merging, and vehicle control. Within the service environment of WATonoBus, the driving behavior decisions are limited to Go/Stop decisions during normal driving and intersection handling, and additional Pullover/Merge, door open/close decisions at bus stop regions. Apart from normal driving decisions, the decision making module will also output emergency stop decisions as the safety fallback strategy. Furthermore, although all the interested obstacles will be reviewed through the decision making module, the final decision will be made with respect to the highest risk obstacle. To achieve the aforementioned objective, finite state machine (FSM) was chosen as the backbone in WATonoBus decision making module (See Fig. 6) because its strengths in structure modularity, interpretation transparency, and robustness to errors or unexpected events [30].
@@
The three important FSM components of the WATonoBus decision making module are states, state transition conditions, and decision-making logic in each state. In an FSM, a state represents a unique situation in the system. In this case, the states are defined based on the static road features in the bus service environment. Specifically, there are normal road segments, all-way stop T-shape intersections, and bus stop regions along the campus ring road. Therefore, the five major states in FSM are defined, namely normal driving, intersection handling, pullover, merging, and passenger on-boarding states. Moreover, to handle the unexpected events, an emergency state is added. A transition in the FSM is a change from one state to another. For WATonoBus decision making module, the transitions are mainly governed by the ego vehicle location and the perception results. Specifically, the ego vehicle location is used to compute the remaining distance to the coming static road features such as all-way stops or bus stop regions. For example, the FSM will change to the pullover state if the ego vehicle is approaching to a bus stop region and then a pullover path will be required. In addition, perception results contain the obstacles around the ego vehicle, indicating whether it is safe to resume bus normal driving. For instance, if there is no vehicle approaching from the rear, the FSM will switch from pullover state to merging state after waiting for passenger on-boarding. Lastly, decisions made under each state can base on different logic due to the traffic rules and vehicle control requirements. The traffic rules in each FSM state list the expected ego vehicle behaviors. To handle an intersection properly, a few traffic rules need to be satisfied in sequence. Assuming there is no pedestrian, the ego vehicle needs to first check if a stop is required for the intersection-waiting vehicle in its lane. It will then be required to stop at the intersection for a while till the intersection is clear. Depending on the arrival order, a decision of whether to yield for the waiting vehicle at other intersection entrances needs to be made. After clearing all these traffic rules check, the ego vehicle can finally start moving and let FSM switch from intersection handling to normal driving state. In addition to traffic rules, the decision-making logic in each state differ from the required vehicle control parameters. For instance, the maximum speed of merging is slower than that of normal driving to ensure the bus to carefully merge out of the bus region while be able to stop immediately for any vehicles suddenly come from the rear (video 1 and 2).
@@ III-D1 Global Path Planning
The global planner determines a high-level path from the current vehicle location to the destination, utilizing HD map information. This algorithm ensures a safe and efficient path, considering any objects or environmental constraints. As WATonoBus operates on a one-lane road, the global planner defines the right curb and center line as the right and left boundaries of the drivable space. The center of this space is assigned as the reference path for the bus to track.
@@ III-D2 Local Path Planning
The task of local path planning is to come up with a more detailed path that follows the global path while taking into account the current state of the vehicle and any obstacles in its immediate surroundings. Moreover, this involves devising a path for the WATonoBus to safely execute pullover and merging maneuvers as needed. These are typically done using potential field [31] and Bezier curves [32]. With the potential field method, we have a mathematical framework to adjust the reference path of the bus whenever obstacles are detected within its drivable space. Following this, smooth segments are generated using Bezier curves from the modified waypoints obtained through the potential field method. For the pullover and merging maneuver, the waypoints are produced using the Bezier curves and control points. By combining the potential field with Bezier curves, the WATonoBus can quickly and safely navigate through its environment and execute its missions.
While following the local path, the objective of the motion controller is to select a desired speed and steering angle for the system to track at each sampling instant. This enables the vehicle to follow the reference path while safely interacting with obstacles (video). The longitudinal controller utilizes a feed-forward algorithm that considers the position and velocity of the ego vehicle relative to obstacles. Firstly, the position $s$ of the WATonoBus and other obstacles along the path is estimated. The distance difference $\Delta s$ between the bus and obstacles along the path is then calculated. If there are no obstacles in the vicinity of the bus, the desired velocity selected by the controller is updated incrementally at each sampling instant using
@@ IV Learning from Daily Operation
We present the handling procedure and lessons learned from both regular bus service and severe weather encountered in WATonoBus daily operations, followed by discussing the insights gained from these case studies.
@@ IV-A Bus Stop Handling
The pullover/merging path planning module ensures a smooth path for the ego vehicle during pullover at bus stops or merging maneuvers (video). Specific waypoints are crucial during pullover, determined considering the bus’s current position, heading, and the identified parking location. Bezier curves are utilized to create a smooth path between these waypoints. Similarly, for merging, waypoints are determined based on the bus’s starting position, orientation, and the target location along its reference path. These waypoints guide the creation of a smooth path using Bezier curves.
(a) Selection
(b) Planning
(c) Boarding
@@ IV-B Intersection Handling
Non-signalized intersections pose challenges due to dynamic traffic flow, right-of-way determination, and potential conflicts (See Fig. 9). Unlike signalized intersections, they lack explicit instructions, demanding advanced perception, DM, and control capabilities for autonomous shuttle buses. Non-signalized intersections rely on right-of-way rules, which can be influenced by factors such as vehicle priority, the order of arrival. Our WATonoBus records the order of arrival for vehicle actors to decide the road priority. When comparing with pedestrian and cyclist at crossings, the shuttle consider itself lowest road priority. Moreover, non-signalized intersections introduce the possibility of conflicts arising from simultaneous movements of different road users. Conflicts can occur due to misjudgment of gaps, ambiguous driver intentions, or violations of right-of-way. The autonomous shuttle bus must be equipped with robust perception systems and prediction modules to detect and track vehicles, pedestrians, and cyclists. To reach as safe and robust behavioral planning for WATonoBus, we use an event-driven technique that reacts to potential conflicts guaranteeing the vehicle is always at safe states.
@@
Figure 9: Intersection handling.
@@ IV-C Perception Performance Under Snow
Our WATonoBus perception system utilizes radar, LiDAR, and camera data fusion, particularly in snowy settings. This integrated approach successfully detects objects in three disparate yet commonplace road scenarios: a van entering the roadway, a car in the opposite lane, and a pedestrian crossing a street, as shown in Fig. 10. For detecting a van entering the road, the radar’s capacity to measure distance and speed plays a crucial role, especially when fused with the camera’s high-definition imaging. The camera contributes valuable details like color and texture, allowing for a highly integrated and predictive object detection framework. This multi-sensory combination anticipates the van’s potential actions, enabling dynamic adjustments for safer autonomous navigation. In the case of a car in the opposite lane, snowy conditions can introduce additional complexities like fog, glare, and reduced contrast. Here, radar’s strong penetration capabilities prove indispensable for initial object detection and range estimation. Combined with the LiDAR’s noisy but detailed spatial data and the camera’s visual cues, the system doesn’t just detect the car; it precisely locates and tracks it, adding another layer of reliability and safety in treacherous weather conditions. Regarding a pedestrian on a crosswalk, snowy conditions often introduce noise into the LiDAR point clouds, making it more challenging to distinguish the shape and position of a pedestrian. In this setting, radar’s Doppler velocity data becomes invaluable. Radar’s robustness to weather interference complements the high-resolution LiDAR data, filtering the noise and ensuring a reliable and comprehensive multi-modal perception of the pedestrian. This fusion mitigates the limitations posed by snowy conditions, providing an enhanced layer of safety and accuracy (video 1 and 2).
@@ V Conclusion
VI Acknowledgement
The authors thank Aaron Sherratt for help with continuous testing and development as well as Jeff Graansma, Adrian Neill, and Michael Duthie. In addition, the authors would like to acknowledge the financial support of the Natural Sciences and Engineering Research Council of Canada, the Canadian Foundation of Innovation, and the Ontario Research Fund.